OpenStudy (anonymous):

Find the set and intrval notation of a range greater than 7

OpenStudy (anonymous):

@geerky42 @Nnesha

OpenStudy (anonymous):

i got y/y<7 for the set

OpenStudy (mathstudent55):

Interval notation uses the smallest and larges value of the interval separated by a comma. Then you use a square bracket to include the value of parentheses to not include the value. With infinity and negative infinity, parentheses are used.

OpenStudy (mathstudent55):

For example, to mean all numbers from 5 to 8, including 5 and 8, you'd write: [5, 8] If it numbers greater than 5 and less than 8, then it's (5, 8) If you want to show the interval from 5 to infinity, then it's \([5, \infty)\) All real numbers is \((- \infty, \infty) \)

OpenStudy (anonymous):

I think I get it now :)

OpenStudy (anonymous):

So, I got y/y>7 for the set notation. Yes?

OpenStudy (mathstudent55):

Set: {the set of all real x, such that x > 7} There are proper math symbols for the above. Interval: \((7, \infty)\)